Solve using any method
(e^3x+1)^2/e^4=e^10x

Answer by stanbon(48535)

(e^(3x)+1)^2=e^(10x+4)
Take the sqrt. of both sides to get:
e^(3x)+1 = e^(5x+2)
I graphed the two sides using a graphing calculator and found
the x-value of the point of intersection to be x=-0.3381103...
Hope this helps.
Cheers,
Stan H.